OpenStudy (anonymous):

If F(x) = x² + 5x and G(x) = 2x + 1, find F(5) + G(6).

OpenStudy (anonymous):

F(x) =( 5² + 5(5))+( 2(6) + 1) = (25+25)+(12+1) =50+13 = 63

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

what about this one If A = {1, 3, 5} and B = {1, 2, 3, 4, 5, 6} then A ∪ B